OpenStudy (anonymous):

Giving Medal! Solve 4m - 2n^2 = 2m - 72 for m a. m = -72 c. m = 72 b. m = 2n^2 - 36 d. m = n^2 - 36

OpenStudy (anonymous):

4m - 2n^2 = 2m - 72 4m - 2m = 2n^2 - 72 m = ?

OpenStudy (anonymous):

well since you didn't put which one you're solving for, I'll put for both: 4m-2n^2=2m-72 subtract both sides by 2m 2m-2n^2= -72 add both sides by 2n^2 2m= 2n^2 -72 divide both sides by 2 m = n^2 - 36 That's for m For n: 4m-2n^2 = 2m - 72 subtract both sides by 4m -2n^2 = -2m -72 multiply both sides by -1 2n^2 = 2m- 72 divide both sides by 2 n^2 = m - 36 take the square root of both sides n = sqrt(m - 36) if m and 36 were being multiplied by each other you can put 6sqrt of m, but since it's addition/subtraction, you can't, it's one term underneath the square root sign